OpenStudy (anonymous):

What is the simplified form of x/y^2 over x^2/y^5? Help please?

OpenStudy (anonymous):

invert and multiply ( actually cancel)

OpenStudy (anonymous):

\[\frac{\frac{a}{b}}{\frac{c}{d}}=\frac{ad}{bc}\]

OpenStudy (anonymous):

is it y^3 over x?

OpenStudy (anonymous):

and so \[\large \frac{\frac{x}{y^2}}{\frac{x^2}{y^5}}\] \[=\frac{xy^5}{y^2x^2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you @satellite73 :)

OpenStudy (anonymous):

yw